OpenStudy (australopithecus):

I have the Series \[\sum_{n=1}^{\infty} (1/e^{n} + 1/n(n+1))\] How do I find the sum of this series as I know it is convergent because if I set 1/e^{n} + 1/n(n+1) = an and take lim an = 0 n->infinity and since the limit of an as n approaches infinity is zero the series is convergent So yeah How do I find the sum of convergent series.

OpenStudy (kinggeorge):

Two things. 1. Just because \[\lim_{n\to\infty}a_n=0\]does not mean that the sum converges. For example, \[\sum_{n=1}^\infty \frac{1}{n}\] 2. split\[\sum_{n=1}^{\infty} (1/e^{n} + 1/n(n+1))=\sum_{n=1}^\infty\frac{1}{e^n}+\sum_{n=1}^\infty \frac{1}{n(n+1)}\]

OpenStudy (anonymous):

first sum is geometric, with \(r=\frac{1}{e}\) second one is going to use partial fractions i bet lets see what KingGeorge says maybe i am wrong

OpenStudy (kinggeorge):

That's exactly what I was going to say ^^

OpenStudy (anonymous):

@Australopithecus do you know how to compute these? first one is the easiest

OpenStudy (kinggeorge):

I have one more thing to add however. Since your sum is from 1 to \(\infty\), you will need to subtract 1 from your geometric sum to account for the \(n=0\) term that you are not including.

OpenStudy (anonymous):

yeah i always make a mistake on that one

OpenStudy (australopithecus):

If the lim an = 0 n->infinity rule does not hold why is it in my text book am I misunderstanding the theory in my text book

OpenStudy (anonymous):

btw lets make sure it is clear why this conveges \[\sum e^{-n}=\sum\frac{1}{e^n}\] converges because it is a geometric series with \(r=\frac{1}{e}<1\)

OpenStudy (anonymous):

and \[\sum\frac{1}{n(n+1)}\] converges because the denominator is a polynomial of degree 2, the numerator is a polynomial of degree 0, and \(2-0=2>1\)

OpenStudy (australopithecus):

right, I understand the first one, but don't you have to check \[\sum_{\infty}^{n-1} 1/n(n+1)\]

OpenStudy (australopithecus):

to determine if the series is convergent?

OpenStudy (australopithecus):

as well

OpenStudy (anonymous):

you check with your eyeballs using the method i wrote above

OpenStudy (anonymous):

if you have a polynomial top and bottom and the denominator is greater in degree by more than one, it converges

OpenStudy (australopithecus):

No I understand that :)

OpenStudy (anonymous):

do you know how to add it?

OpenStudy (australopithecus):

yes -_-

OpenStudy (anonymous):

ok then you are in good shape right?

OpenStudy (australopithecus):

well I'm just confused by the theorem that involves taking the limit of an

OpenStudy (australopithecus):

Oh I get it now

OpenStudy (kinggeorge):

I think you're confusing it with the theorem that says: If \(\lim_{n\to\infty} a_n \neq0\) then the series diverges. Note that this only tells you if it diverges, and nothing about convergence.

OpenStudy (australopithecus):

if we find that the limit of an = 0 then we know nothing about convergence or divergence but if we find anything else as the limit we know it is divergent

OpenStudy (kinggeorge):

Precisely.